OpenStudy (anonymous):

Given two linear equations, explain how you can determine if the lines are parallel without graphing them.

OpenStudy (anonymous):

What are "given 2 linear equations"?

OpenStudy (anonymous):

any two l. equations

OpenStudy (anonymous):

use this eqto compare y=mx+c see if value of m is same in both eq.then thay are paralell.. where m=slope

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

I give you the way to see whether they are // or not First off, no matter what the form is, both equations MUST have the same form If you have the forms of y = \(\color{red}{m}\)x +b and If your equations have the same \(\color{red}{m}\), then they are // For example : y = \(\color{red}{5}\) x+7 y = \(\color{red}{5}\) x+9 then, they are parallel If you have the forms of ay + bx =c, then you must have a, b are the same on both equations, then they are parallel For example: \(\color{red}{5}\) y+\(\color{red}{7}\) x=7 \(\color{red}{5}\) y+\(\color{red}{7}\) x=15 then, they are parallel If you have the first equation has the form y = mx +b and the second is in the form of ax +by =c you Must convert the first or the second one to have them be in the same form to consider.